Patricia is considering investing in a company's stock and is aware that the return on that investment is particularly sensitive to how the economy is performing. Her analysis suggests that four states of the economy can affect the return on the investment.
Probability         Return
Boom 0.4 25.00%
Good 0.1 15.00%
Level 0.2 10.00%
Slump 0.3 -5.00%
Use the table of returns and probabilities above to determine the expected return on Patricia’s investment? (Round answer to 3 decimal places, e.g. 0.076.)
Expected return

Use the table of returns and probabilities above to determine the standard deviation of the return on Patricia's investment? (Round answer to 5 decimal places, e.g. 0.07680.)
Standard deviation
Answer #1

Expected return of portfolio E(r) =  \sum_{}^{} p(s)*r(s),

where p(s) is the probability of each scenario,

and r(s) is the expected return of each scenario.

Variance of portfolio \sigma 2  = \sum_{}^{} p(s)*[r(s) - E(r)]2

where [r(s) - E(r)]2 is the squared deviation from the expected return.

Standard deviation =  \sqrt{}variance

Expected return = 12.000%

Standard deviation = 0.01560

Expected Probability return r(s) p(s) * ((r(s) - E(r))^2) 1 Scenariop(s) P(s) * r(s) 25.00% 2 Boom 0.4 0.1 0.00676 3 Good 15.

Probability p(s) p(s) * ((r(s) - E(r))^2) Expected return r(s) p(s) * r(s) 1 Scenario =B2*((C2-$D$6)^2) =B3*((C3-$D$6)^2) =B4